40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10034     Entry time: Thu Jun 12 16:56:31 2014
Author: Nichin 
Type: Update 
Category: Electronics 
Subject: PD Inspection 

I and Eric Gustafson inspected the automated PD frequency response measurement system which Alex Cole built last summer. We just lifted the tops off the tables [AS table, POY table and ITMX table] and looked at the alignment checking to see if the correct optical fibers from the fiber splitter were illuminating the correct photodiodes. We did not change anything at all and put the covers back on the tables.

The PDF attached shows the state of each PD fiber pair.  The fibers labeled REFL11 and REFL55 were reversed and illuminating the wrong photodiodes.

We will do a manual measurement of REFL33 tomorrow using the network analyzer and the modulatable laser but not the RF switch.  Afterward we will check to make sure the RF cables are connected to the correct channels of the RF switch according to the switch list (/users/alex.cole/switchList).

Attachment 1: Inspection_PD_Freq_Resonse_system_12th_June_2014.pdf  351 kB  | Hide | Hide all
Inspection_PD_Freq_Resonse_system_12th_June_2014.pdf Inspection_PD_Freq_Resonse_system_12th_June_2014.pdf
ELOG V3.1.3-